A sphere of radius r starts from rest and rolls without slipping along a curved surface, dropping through a vertical distance of 0.431 m. Find the final speed v of the sphere's center of mass.
Answer #1

Energy conservation

Potential energy = rolling kinetic energy + translation kinetic energy

mgh= 1/2mv^2 + 1/2 I *w^2

I=2/5 *mr^2 , w=v/r

0.7v^2= 9.81*0.431, v= 2.46 m/s